# Explore the properties of $\Lambda(1670)$ in the Cabibbo-favored process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay

### v1

@@ Abstract
Recently, the Belle and LHCb Collaborations have measured the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay and reported the $pK^{-}$ invariant mass distribution, which shows a clear cusp structure around the $\eta\Lambda$ threshold. In this work, we have analyzed this process by considering the triangle mechanism and the $S$-wave pseudoscalar meson-octet baryon interactions within the chiral unitary approach, which dynamically generate the $\Lambda(1670)$. Our results are in good agreement with the Belle measurements, which implies that the cusp structure around $\eta\Lambda$ threshold could be associated with the $\Lambda(1670)$ with the molecular nature.
@@ I Introduction
Since the discovery of the charmed baryon $\Lambda_{c}^{+}$ Knapp:1976qw , there has been many studies of the hadronic weak decays of charmed baryons, which provide an important platform to study the strong and weak interactions. Yet, both the experimental and theoretical progresses in the study of hadronic decays of charmed baryons were very slow for a long time. This situation was reversed since 2014 as there were several major breakthroughs in charmed-baryon experiments Cheng:2021qpd ; Zhong:2024zme . In the last decade, there are a lot of multi-body hadronic decay modes of the charmed baryons reported by the BESIII, Belle, and LHCb Collaborations ParticleDataGroup:2022pth ; BESIII:2018qyg ; BESIII:2023sdr ; Belle:2022pwd ; LHCb:2017xtf ; LHCb:2019ldj , which involve complicated final state interactions, and have motivated many studies of the light hadrons in those processes, such as the scalar mesons $a_{0}(980)/f_{0}(980)$ in $\Lambda^{+}_{c}\to pK^{+}K^{-},p\pi^{+}\pi^{-}$, $\Lambda^{+}_{c}\to p\eta\pi^{+}$ Feng:2020jvp , and light baryons in $\Lambda^{+}_{c}\to{\pi}^{0}{\phi}p$ Xie:2017mbe , $\Lambda^{+}_{c}\to\bar{K^{0}}\eta p$ Xie:2017erh , $\Lambda^{+}_{c}\to\Lambda\eta\pi^{+}$ Wang:2022nac ; Xie:2017xwx ; Xie:2016evi ; Lyu:2024qgc , $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ Liu:2019dqc .
In 2023, the LHCb Collaboration has performed an amplitude analysis of the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay in connection with the semi-leptonic beauty hadron decays LHCb:2022sck , where a cusp structure around $\eta\Lambda$ threshold appears in the measured $pK^{-}$ invariant mass distribution. Later, the Belle collaboration has also measured the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$, and found a narrow peaking structure near the $\Lambda\eta$ threshold in the $pK^{-}$ invariant-mass distribution Belle:2022cbs . The Belle has explained this structure within the two approaches, the Breit-Wigner function of a new resonance, and a Flatté function, and suggested that the cusp structure around the $\eta\Lambda$ could be associated with $\Lambda(1670)$ Belle:2022cbs . Indeed, the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ has been studied for a long time, and its complex resonant structure has attracted widespread attention from theorist and experimentalist Filaseta:1987rx ; BESIII:2015bjk ; Liu:2019dqc ; Miyahara:2015cja .
@@
As we know, the ground octet and decuplet baryons could be well described within the naive quark model, however the low-lying excited baryons with the quantum numbers of spin-parity $J^{P}=1/2^{-}$ have some exotic properties, which are difficult to be explained within the naive quark model. For instance, the $\Lambda(1405)$ mass is difficult to understand in the naive quark model as a $P$-wave excited baryon Isgur:1978xj , and this state could be generated by the $\bar{K}N$ coupled-channel interaction within the unitary chiral approach Kaiser:1995eg ; Oset:1997it ; Oller:2000fj .
Concerning the low-lying excited baryon $\Lambda(1670)$, it exhibits itself as an enhancement structure in the $\eta\Lambda$ invariant mass distribution of the processes $K^{-}p\to\eta\Lambda$ CrystalBall:2001uhc and $\Lambda_{c}^{+}\to\eta\Lambda\pi^{+}$ BESIII:2018qyg ; Belle:2020xku , and also a dip structure around $\eta\Lambda$ threshold in the $\bar{K}^{0}n$ invariant mass distribution of the process $K^{-}p\to\bar{K}^{0}n$ Gopal:1976gs . Indeed, the dip structure could be produced by the interference between the resonance and the other contributions Feng:2020jvp ; Guo:2019twa ; Ding:2023eps ; Lu:2016roh ; Chen:2015sxa , and one nice example is that the $f_{0}(980)$ appears as a dip for the process $J/\psi\to\omega\pi^{+}\pi^{-}$ BES:2004mws , but as a sharp peak in the $\pi^{+}\pi^{-}$ invariant mass distribution for the process $J/\psi\to\phi\pi^{+}\pi^{-}$ BES:2004twe .
@@ II Formalism
The process $\Lambda_{c}^{+}\to a_{0}(980)^{+}\Lambda$ could happen via the following steps. First, the $c$ quark of the initial $\Lambda^{+}_{c}$ weakly decays into an $s$ quark plus a $W^{+}$ boson, then the $W^{+}$ boson goes into a $u\bar{d}$ pair, as shown in Fig. 1. Second, the $u\bar{d}$ pair from the $W^{+}$ decay will hadronize into $a_{0}(980)^{+}$, while the $s$ quark from the $c$ decay, and the $ud$ pair of the initial $\Lambda^{+}_{c}$, combine to give the state $\Lambda$.
@@
Figure 2: Triangle mechanism for the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$. Kinematic conventions for the intermediate states are $\Lambda_{c}^{+}(P)$, $\Lambda(q_{1})$, $a_{0}(980)(P-q)$, $\pi^{+}(k)$ and $\eta(P-q-k)$.
Firstly, we consider the decay of $\Lambda_{c}\to\Lambda a_{0}(980)$ via $S$-wave, and the amplitude could be given as,
@@
where the constant $g_{\Lambda_{c}^{+}a_{0}\Lambda}$ corresponds to the effective coupling between $\Lambda_{c}^{+}$ and $a_{0}(980)^{+}\Lambda$. And the amplitude for the vertex of $a_{0}(980)^{+}\to\eta\pi$ can be written as,
@@
with the effective coupling $g_{a_{0}\eta\pi}$ between $a_{0}(980)^{+}$ and $\eta\pi$.
@@
where $m_{l}$ and $M_{l}$ are the masses of meson and baryon of the $l$-th channel. Here, we take $\mu=630$ MeV. Since the pole position of $\Lambda(1670)$ is quite sensitive to the value of $a_{K\Xi}$ and only moderately sensitive to $a_{\bar{K}N}$, $a_{\pi\Sigma}$, $a_{\eta\Lambda}$, we take $a_{K\Xi}$ to be a free parameter, and adopt the values $a_{\bar{K}N}=-1.84$, $a_{\pi\Sigma}=-2.00$, and $a_{\eta\Lambda}=-2.25$ from Ref. Oset:2001cn ; Wang:2022nac ; Lyu:2024qgc .
@@
with $P^{0}=M_{\Lambda_{c}^{+}}$, $\omega_{\Lambda}=\sqrt{|\vec{q}\,|^{2}+m_{\Lambda}^{2}}$, $\omega_{\eta}=\sqrt{|\vec{q}+\vec{k}|^{2}+m_{\eta}^{2}}$, and $\omega_{a_{0}}=\sqrt{|\vec{q}\,|^{2}+m_{a_{0}}^{2}}$. Here $m_{a_{0}}=980$ MeV, and $\Gamma_{a_{0}}=75$ MeVParticleDataGroup:2020ssz ; Liu:2019dqc . $k^{0}$ and $\vec{k}$ are the energy and three-momentum of the $\pi^{+}$ emitted from $a_{0}(980)$ in the rest frame of the $\Lambda_{c}^{+}$, which are given by
@@
In this work, to regularize the integral of Eq. (9), we use the cutoff $q_{\rm max}=1000$ MeV.
@@
$\frac{d\Gamma}{dM_{K^{-}p}}=\frac{1}{(2\pi)^{3}}\frac{p_{\pi^{+}}\tilde{p}_{K^{-}}}{4M^{2}_{\Lambda_{c}^{+}}}|\mathcal{M}^{\rm TS}|^{2},$ (12)
@@
with Källen function $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2zx$. All the masses of the mesons and baryons involved in our calculations are taken from Review of Particle Physics (RPP) ParticleDataGroup:2022pth .
@@ III Numerical results and discussion
Figure 3: (Color online) Invariant $pK^{-}$ mass distribution for the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay considering the triangle mechanism.
@@
$\mathcal{Q}^{\prime}$ $(0.178\pm 0.002)\times 10^{9}$
@@
Figure 4: (Color online) Invariant $pK^{-}$ mass distribution for the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay considering the TS mechanism with fited parameters.
@@
One can find that, our results are in agreement with the imaginary part of the $\bar{K}N\to\bar{K}N$ transition amplitude within uncertainties, while our prediction for the real part below the $\eta\Lambda$ threshold differs from the data Gopal:1976gs , which implies that a large constant background contribution should be taken into account Oset:2001cn . Our results of the $\bar{K}N\to\pi\Sigma$ transition amplitude show the resonant feature with the same pattern as the experiment Gopal:1976gs . The disagreement of the imaginary part with the data also implies an apparent background contribution missed in our model Oset:2001cn .
Figure 6: (Color online) According to the calculation results in Ref.Oset:2001cn , $a_{K\Xi}$ takes the amplitude of -2.776 in the $I=0$ channel. Experimental data are taken from Ref. Gopal:1976gs . (a) and (b) are the real and imaginary parts of $\bar{K}N\to\bar{K}N$ amplitude, (c) and (d) are the real and imaginary parts of $\bar{K}N\to\pi\Sigma$ amplitude.
@@ IV Summary
We have shown that the pole position of the $\Lambda(1670)$ is quite sensitive to the subtract constant $a_{K\Xi}$ that regularizes the meson-baryon loop functions. By fitting to the $K^{-}p$ invariant mass distribution of the efficiency-corrected events measured by Belle, we have obtained the $\chi^{2}/N_{d.o.f.}=1.25$, and the fitted subtract constant $a_{K\Xi}=-2.776\pm 0.003$, which corresponds to the pole of $(1669.2,20.4i)$ MeV for the $\Lambda(1670)$. Our results are in good agreement with the $K^{-}p$ invariant mass distribution of the Belle data, which supports that the observed cusp structure around $\eta\Lambda$ threshold could be associated with the $\Lambda(1670)$.
@@ V ACKNOWLEDGMENTS
We would like to thank Eulogio Oset for useful comments. This work is supported by the Natural Science Foundation of Henan under Grant No. 232300421140 and No. 222300420554, the National Natural Science Foundation of China under Grant No. 12192263, the Project of Youth Backbone Teachers of Colleges and Universities of Henan Province (2020GGJS017), and the Open Project of Guangxi Key Laboratory of Nuclear Physics and Nuclear Technology, No. NLK2021-08.

### v2

@@ Abstract
Recently, the Belle and LHCb Collaborations have measured the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay and reported the $K^{-}p$ invariant mass distribution, which shows a clear cusp structure around the $\eta\Lambda$ threshold. In this work, we have analyzed this process by considering the triangle mechanism and the $S$-wave pseudoscalar meson-octet baryon interactions within the chiral unitary approach, which dynamically generates the $\Lambda(1670)$. Our results are in good agreement with the Belle measurements, which implies that the cusp structure around $\eta\Lambda$ threshold could be associated with the $\Lambda(1670)$ with the molecular nature.
@@ I Introduction
Since the discovery of the charmed baryon $\Lambda_{c}^{+}$ Knapp:1976qw , there has been many studies of the hadronic weak decays of charmed baryons, which provide an important platform to study the strong and weak interactions Miyahara:2015cja ; Jia:2024pyb ; Geng:2018upx ; Geng:2024sgq . Yet, both the experimental and theoretical progresses in the study of hadronic decays of charmed baryons were very slow for a long time. This situation was reversed since 2014 as there were several major breakthroughs in charmed-baryon experiments Cheng:2021qpd ; Zhong:2024zme . In the last decade, there are a lot of multi-body hadronic decay modes of the charmed baryons reported by the BESIII, Belle, and LHCb Collaborations ParticleDataGroup:2024cfk ; BESIII:2018qyg ; BESIII:2023sdr ; Belle:2022pwd ; LHCb:2017xtf ; LHCb:2019ldj , which involve complicated final state interactions, and have motivated many studies of the light hadrons in those processes, such as the scalar mesons $a_{0}(980)/f_{0}(980)$ in $\Lambda^{+}_{c}\to pK^{+}K^{-},p\pi^{+}\pi^{-}$ Wang:2020pem , $\Lambda^{+}_{c}\to p\eta\pi^{+}$ Feng:2020jvp , and light baryons in $\Lambda^{+}_{c}\to{\pi}^{0}{\phi}p$ Xie:2017mbe , $\Lambda^{+}_{c}\to\bar{K}^{0}\eta p$ Xie:2017erh ; Li:2024rqb , $\Lambda^{+}_{c}\to\Lambda\eta\pi^{+}$ Wang:2022nac ; Xie:2017xwx ; Xie:2016evi ; Lyu:2024qgc , $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ Liu:2019dqc .
In 2023, the LHCb Collaboration has performed an amplitude analysis of the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay in connection with the semi-leptonic beauty hadron decays LHCb:2022sck , where a cusp structure around $\eta\Lambda$ threshold appears in the measured $K^{-}p$ invariant mass distribution. Later, the Belle Collaboration has also measured the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$, and found a narrow peaking structure near the $\Lambda\eta$ threshold in the $K^{-}p$ invariant-mass distribution Belle:2022cbs . The Belle has explained this structure within the two approaches, the Breit-Wigner function of a new resonance, and a Flatté function, and suggested that the cusp structure around the $\eta\Lambda$ could be associated with $\Lambda(1670)$ Belle:2022cbs . Indeed, the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ has been studied for a long time, and its complex resonant structure has attracted widespread attention from theorists and experimentalists Filaseta:1987rx ; BESIII:2015bjk ; Liu:2019dqc ; Miyahara:2015cja ; Duan:2024okk .
@@
As we know, the ground octet and decuplet baryons could be well described within the naive quark model, however the low-lying excited baryons with the quantum numbers of spin-parity $J^{P}=1/2^{-}$ have some exotic properties, which are difficult to be explained within the naive quark model. For instance Wang:2024jyk , the $\Lambda(1405)$ mass is difficult to understand in the naive quark model as a $P$-wave excited baryon Isgur:1978xj , and this state could be generated by the $\bar{K}N$ coupled-channel interaction within the unitary chiral approach Kaiser:1995eg ; Oset:1997it ; Oller:2000fj . Concerning the low-lying excited baryon $\Lambda(1670)$, it exhibits itself as an enhancement structure in the $\eta\Lambda$ invariant mass distribution of the processes $K^{-}p\to\eta\Lambda$ CrystalBall:2001uhc and $\Lambda_{c}^{+}\to\eta\Lambda\pi^{+}$ BESIII:2018qyg ; Belle:2020xku , and also a dip structure around $\eta\Lambda$ threshold in the $\bar{K}^{0}n$ invariant mass distribution of the process $K^{-}p\to\bar{K}^{0}n$ Gopal:1976gs . Indeed, the dip structure could be produced by the interference between the resonance and the other contributions Feng:2020jvp ; Guo:2019twa ; Ding:2023eps ; Lu:2016roh ; Chen:2015sxa , and one nice example is that the $f_{0}(980)$ appears as a dip for the process $J/\psi\to\omega\pi^{+}\pi^{-}$ BES:2004mws , but as a sharp peak in the $\pi^{+}\pi^{-}$ invariant mass distribution for the process $J/\psi\to\phi\pi^{+}\pi^{-}$ BES:2004twe .
@@
In Ref. Menadue:2011pd , the authors have performed the first investigation of the low-lying $J^{P}=1/2^{-}$ spectrum of the $\Lambda$ baryon at near-physical pion masses, and identified a low-lying $\Lambda(1405)$ using conventional three-quark operators. Meanwhile, two $\Lambda$ states are found in the $\Lambda(1670)$ mass region, which implies that a single particle state makes an important role to the structure of the $\Lambda(1670)$.
@@ II Formalism
The process $\Lambda_{c}^{+}\to a_{0}(980)^{+}\Lambda$ could happen via the following steps. First, the $c$ quark of the initial $\Lambda^{+}_{c}$ weakly decays into an $s$ quark plus a $W^{+}$ boson, then the $W^{+}$ boson goes into a $u\bar{d}$ pair, as shown in Fig. 1. Second, the $u\bar{d}$ pair from the $W^{+}$ decay will hadronize into $a_{0}(980)^{+}$, while the $s$ quark from the $c$ decay, and the $ud$ pair of the initial $\Lambda^{+}_{c}$, combine to give the state $\Lambda$ Wang:2022nac ; Lyu:2024qgc .
@@
Figure 2: Triangle mechanism for the process $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$. Kinematic conventions for the intermediate states are $\Lambda_{c}^{+}(P)$, $\Lambda(q)$, $a_{0}(980)(P-q)$, $\pi^{+}(k)$ and $\eta(P-q-k)$.
Firstly, we consider the decay of $\Lambda_{c}^{+}\to\Lambda a_{0}(980)$ via $S$-wave, and the amplitude could be given as,
@@
where the constant $g_{\Lambda_{c}^{+}a_{0}\Lambda}$ corresponds to the effective coupling between $\Lambda_{c}^{+}$ and $a_{0}(980)^{+}\Lambda$. And the amplitude for the vertex of $a_{0}(980)^{+}\to\eta\pi^{+}$ can be written as,
@@
with the effective coupling $g_{a_{0}\eta\pi}$ between $a_{0}(980)^{+}$ and $\eta\pi^{+}$.
@@
where $m_{l}$ and $M_{l}$ are the masses of meson and baryon of the $l$-th channel. Here, we take $\mu=630$ MeV [^1] The changes in the $\mu$ can be accommodated in terms of changes in the subtraction constant $a_{l}$ Oset:2001cn .. Since the pole position of $\Lambda(1670)$ is quite sensitive to the value of $a_{K\Xi}$ and only moderately sensitive to the subtraction constant $a_{\bar{K}N}$, $a_{\pi\Sigma}$, $a_{\eta\Lambda}$, we take $a_{K\Xi}$ to be a free parameter, and adopt the values $a_{\bar{K}N}=-1.84$, $a_{\pi\Sigma}=-2.00$, and $a_{\eta\Lambda}=-2.25$ from Refs. Oset:2001cn ; Wang:2022nac ; Lyu:2024qgc .
@@
with $P^{0}=M_{\Lambda_{c}^{+}}$, $\omega_{\Lambda}=\sqrt{|\vec{q}\,|^{2}+m_{\Lambda}^{2}}$, $\omega_{\eta}=\sqrt{|\vec{q}+\vec{k}|^{2}+m_{\eta}^{2}}$, and $\omega_{a_{0}}=\sqrt{|\vec{q}\,|^{2}+m_{a_{0}}^{2}}$. Here $m_{a_{0}}=980$ MeV, and $\Gamma_{a_{0}}=75$ MeV ParticleDataGroup:2024cfk ; Liu:2019dqc  [^2] As we know, the width of the scalar $a_{0}(980)$ has not been well established due to its exotic structure, thus we take the centre value of the estimated range $50\sim 100$ MeV of Particle Data Group (PDG) ParticleDataGroup:2024cfk . Indeed, it has been shown in Ref. Liu:2019dqc that the width of $a_{0}(980)$ has less influence on the contribution from the triangle mechanism.. $k^{0}$ and $\vec{k}$ are the energy and three-momentum of the $\pi^{+}$ emitted from $a_{0}(980)$ in the rest frame of the $\Lambda_{c}^{+}$, which are given by
@@
In this work, to regularize the integral of Eq. (9), we use the cutoff $q_{\rm max}=1000$ MeV, and will show that our results are insensitive to its value.
@@
$\frac{d\Gamma}{dM_{K^{-}p}}=\frac{1}{(2\pi)^{3}}\frac{2M_{p}2M_{\Lambda_{c}}p_{\pi^{+}}\tilde{p}_{K^{-}}}{4M^{2}_{\Lambda_{c}^{+}}}|{\mathcal{T}^{\rm TS}}|^{2},$ (12)
@@
with Källen function $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2yz-2zx$. All the masses of the mesons and baryons involved in our calculations are taken from PDG ParticleDataGroup:2024cfk .
@@ III Numerical results and discussion
Figure 3: (Color online) Invariant $K^{-}p$ mass distribution for the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay considering the triangle mechanism.
Figure 4: (Color online) Invariant $K^{-}p$ mass distribution for the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay considering the triangle mechanism with $q_{\rm max}=800,1000,1200$ MeV.
@@
$\mathcal{Q}^{\prime}$ $(6.092\pm 0.068)\times 10^{4}$
@@
Figure 5: (Color online) Invariant $K^{-}p$ mass distribution for the $\Lambda^{+}_{c}\to pK^{-}\pi^{+}$ decay considering the triangle mechanism with fitted parameters.
@@
Meanwhile, in Fig. 4 we also show the results with $a_{K\Xi}=-2.80$ and different values of the cut-off of Eq. (9), i.e. $q_{\rm max}=800,1000,1200$ MeV, and one can find that the line shape and the cusp position have no sizeable changes. Thus, in the following we set $q_{\rm max}=1000$ MeV.
@@
One can find that, our results are in agreement with the imaginary part of the $\bar{K}N\to\bar{K}N$ transition amplitude within uncertainties, while our prediction for the real part below the $\eta\Lambda$ threshold differs from the data Gopal:1976gs , which implies that a large constant background contribution should be taken into account Oset:2001cn . Our results of the $\bar{K}N\to\pi\Sigma$ transition amplitude show the resonant feature with the same pattern as the experiment Gopal:1976gs . The disagreement of the imaginary part with the data also implies an apparent background contributions missed in our model Oset:2001cn , such as the $t/u$-channel exchange terms.
Figure 7: (Color online)The figures (a) and (b) are the real part and imaginary part of the $\bar{K}N\to\bar{K}N$ amplitude, while the figures (c) and (d) are the real part and imaginary part of $\bar{K}N\to\pi\Sigma$ amplitude. Here $a_{K\Xi}=-2.776$ and the experimental data are taken from Ref. Gopal:1976gs .
@@ IV Summary
We have shown that the pole position of the $\Lambda(1670)$ is quite sensitive to the subtract constant $a_{K\Xi}$ that regularizes the meson-baryon loop functions. By fitting to the $K^{-}p$ invariant mass distribution of the efficiency-corrected events measured by Belle, we have obtained the $\chi^{2}/N_{d.o.f.}=1.25$, and the fitted subtract constant $a_{K\Xi}=-2.776\pm 0.003$, which corresponds to the pole of $(1669.0,20.3i)$ MeV for the $\Lambda(1670)$. Our results are in good agreement with the $K^{-}p$ invariant mass distribution of the Belle data, which supports that the observed cusp structure around $\eta\Lambda$ threshold could be associated with the $\Lambda(1670)$.
@@
Recently, Ref. Liu:2023xvy has studied two different scenarios for the internal structure of the $\Lambda(1670)$. One scenario assumes that the $\Lambda(1670)$ is dynamically generated through meson-baryon scattering, and the other assumes that the $\Lambda(1670)$ is a bare quark-model-like state mixing with these $I=0$ interacting channels. It is concluded that one cannot distinguish those two scenarios using scattering data alone, and exploring the structure of the $\Lambda(1670)$ in the finite volume of lattice QCD is necessary Liu:2023xvy . In addition, recent lattice QCD results BaryonScatteringBaSc:2023zvt ; BaryonScatteringBaSc:2023ori for the $J^{P}=1/2^{-}$ $\Lambda$ channel implies that, including a single particle contribution in the $\Lambda(1670)$ structure is important. Thus, the internal structure of the $\Lambda(1670)$ is not well understood, and the more precise measurements could shed light on this question.
@@ V ACKNOWLEDGMENTS
We would like to thank Eulogio Oset for useful comments. This work is supported by the National Key R&D Program of China (No. 2024YFE0105200), the Natural Science Foundation of Henan under Grant No. 232300421140 and No. 222300420554, the National Natural Science Foundation of China under Grant No. 12205075, No. 12475086, and No. 12192263.